OpenStudy (mehek14):

The price of admission o the New Haven Children's Museum is $6 more for adult than for children. If a group of 6 adults and 18 children pays a total of $108 to visit the museum, what is the cost, in dollars, of the ticket for one adult? (A) $7.00 (B) $9.00 (C) $10.00 (D) $12.00 (E) $13.00

Mehek (mehek14):

@DanJS

Mehek (mehek14):

@Nnesha

OpenStudy (danjs):

The problem gives you information to make 2 equations.

OpenStudy (danjs):

$6 more for adult than for children Cost of Adult = Cost of Child + 6 A = C + 6

OpenStudy (danjs):

6 adults and 18 children pays a total of $108 6*(cost of adult) + 18*(cost of child) = 108

OpenStudy (danjs):

2 equations, 2 unknowns A = C + 6 6A + 18C = 108

OpenStudy (danjs):

Sub, the first into the second, and solve for C.

OpenStudy (danjs):

Or since they want just the cost for an adult. A. Solve the first for C, then put C = A - 6, into the second, and solve for A

Mehek (mehek14):

I got c = 3 so since it's 6 more for adults it's going to be ( so B?

Mehek (mehek14):

@DanJS

OpenStudy (danjs):

6C+36+18C=108 24C = 72 C=3 A=C+6 = 9

Mehek (mehek14):

thanks :)